OpenStudy (anonymous):

Batteries of each type have lifetimes before recharge that can be assumed independent and Normally distributed. The mean and standard deviation of the lifetimes of the Duxcell batteries are 10 and 2 minutes respectively, the mean and standard deviation for the Infinitycell batteries are 14 and 3 minutes respectively. Part a) What is the probability that a Duxcell battery will last longer than an Infinitycell battery?

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

So i did P(D>14) so 1 - P(Z < 2) = 1 - 0.9772 = 0.0228. where did i go wrong?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

you completely ignored the standard deviations

OpenStudy (amistre64):

you might want to perform a hyptothesis testing is one idea